if you add5 to the sum of bubba's age and his poppa's age the total is 48. bubba's poppa is 5 years less than 3 times the age of bubba. how old is bubba and his poppa?
Answer by josmiceli(19441)

b + p + 5 = 48
p = 3b - 5
b + 3b -5 + 5 = 48
4b = 48
b = 12
p = 3*(12) - 5
p = 36 - 5
p = 31
so, bubba is 12 and poppa is 31
check
12 + 31 + 5 = 48
48 = 48
OK